Create a proposal to start a new company. You will sell one product only and will need to have a variety of fixed and variable costs in the make-shift company. Fictionalize a data set and use this data to generate a CVP analysis showing both the break-even point and also the number of units to be sold in order to earn a $10,000 profit. Sale price must be 200 Please list 5 variable costs and 5 fixed costs
Answer #1

Sale price = $200

Target profit = $10,000

Let variable cost be $20, $40, $50, $80 and $100 and fixed costs be $3,600, $8,000, $15,000, $24,000 and $50,000 respectively.

Contribution margin per unit = Sale price - Variable cost

Break even point = Fixed cost/Contribution margin per unit

Number of units to be sold in order to earn a target profit = (Fixed cost + Target profit)/Contribution margin per unit

Variable cost $20, Fixed cost $3,600 Variable cost $40, Fixed cost $8,000 Variable cost $50, Fixed cost $15,000 Variable cost $80, Fixed cost $24,000 Variable cost $100, Fixed cost $50,000
Contribution margin per unit $180 $160 $150 $120 $100
Break even point

3,600/180 = 20 units

8,000/160 = 50 units 15,000/150 = 100 units 24,000/120 = 200 units 50,000/100 = 500 units
Number of units to be sold in order to earn a $10,000 profit

(3,600 + 10,000)/180

13,600/180 = 76 units

(8,000 + 10,000)/160

= 18,000/160 = 113 units

(15,000 + 10,000)/150 = 25,000/150

= 167 units

(24,000 + 10,000)/120 = 34,000/120

= 283 units

(50,000 + 10,000)/100 = 60,000/100

= 600 units

Kindly give a positive rating if you are satisfied with the answer. Feel free to ask if you have any doubts. Thanks.